if p = 2l + 2w what is the perimeter of a rectangle that is 6 in x 7 in ? could you please explain to me how to do this, because my teacher does not...
Answer by rapaljer(4551)

It might help to draw a picture of a rectangle, and label the four sides. The perimeter is actually just what you get when you add up the four sides.

P= 6+7+6+7= 26 in.

Now, as a shortcut for this, notice that when you added up the four sides, you actually were adding TWO widths and TWO lengths. This is how we get that famous fomula for perimeter of a rectangle:

P= 2W + 2L

So, the other way to do this:
P= 2*6 + 2*7
P = 12 + 14 = 26 in.

R^2 at SCC